# Variational Calculation of the Hyperfine Stark Effect in Atomic 87Rb, 133Cs, and 169Tm

### v1

@@ III.1.4 Finite-Field (FF) Method
For the electronic ground state corresponding to the valence configuration $4s^{1}$ the total electronic angular momentum quantum number is ${J=1/2}$. The considered isotope has $I=3/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=2$ and $F_{l}=1$. The fractional occupation in the DCHF calculation is $f=1/2$.
Results for the hyperfine Stark coefficient are shown in Table 1 for various electronic-structure models and are compared with experimental and theoretical literature results. The result in mean-field approximation (DCHF) differs from the experimental result by about $20$% and lowest-order electron correlation effects from the $3s,3p,4s$ Rb shells even increase this discrepancy (model SD9). Including virtual spinors up to $30$ a.u.  does not lead to a significant change, so for the computationally most expensive models the virtual cutoff is set to $10$ a.u.
Upon including combined triple excitations from the $3s,3p$ shells and the $4s$ valence shell (model SD8_SDT9) a strong correction to $k$ is obtained that, however, seems to overshoot the exact value. The inclusion of full triple excitations (model SDT9), i.e., including those from the $3s,3p$ shells, proves to be crucial in obtaining an accurate value for $k$. Excellent agreement with the result obtained by Safronova et al. Safronova_BBR_2009 is achieved at this level of calculation.
Table 1: Stark hyperfine coefficient $k$ for 87Rb with nuclear magnetic moment $\mu=2.75131\mu_{N}$ ISOLDE_BWeffect_1993 .
Model $k$ [$10^{-10}$ Hz/((V/m)2)]
DCHF $-1.546$
SD9/10au $-1.662$
SD9/30au $-1.669$
SD8_SDT9/30au $-1.202$
SDT9/10au $-1.265$
SDT9/30au $-1.271$
Exp. Mowat_Li_1972 $-1.23(3)$
Safronova et al. Safronova_BBR_2009 $-1.272$
@@
As a general effect, the hyperfine constant of a given atom in an electronic $S$ state diminishes when an external $E$ field is included. This observation is explained by the fact that the $E$ field partially shifts spin density from $s$ wave to $p$ wave character in the atomic ground state (through $s$-$p$ mixing in the polarized atom), thus reducing the hyperfine interaction. On the other hand, electron correlation effects increase the hyperfine constants of $S$ state atoms, which means that effectively spin density residing in $s$ spinors is increased by allowing for the mixing in of determinants with open-shell character.
@@
This important isotope of cesium serves as a second test case. For the electronic ground state corresponding to the valence configuration $5s^{1}$ the total electronic angular momentum quantum number is ${J=1/2}$. The considered isotope has $I=7/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=4$ and $F_{l}=3$. The fractional occupation in the DCHF calculation is $f=1/2$. Results for 133Cs are compiled in Table 2.
@@
As in the case of 87Rb the mean-field result for $k$ is too large on the absolute, compared with the cited reference values, by over $40$%. Again, the inclusion of lowest-order electron correlation effects even increases this deviation. It is here in addition and for the model SD9 shown that the consideration of virtual spinors of very high energy (up to $1000$ a.u. ) in the wave-function expansion does not affect the results significantly. Also as for 87Rb combined triple excitations from the outermost shells yield the decisive correction to $k$. Including full triple excitations (model SDT9/30au) yields a result that deviates from the most accurate literature results by roughly $+9$%. Accounting for combined quadruple excitations quenches the deviation to $-7$% where again, as in the Rb atom, the correction slightly overshoots the exact value. This latter model comprises about $1.9$ billion ($10^{9}$) terms in the CI expansion.
@@
Table 2: Stark hyperfine frequency shift $k$ for 133Cs ($5s_{1/2}$), $I=3.5$, $E_{z}=0.001$ a.u., hyperfine level quantum numbers are $F_{u}=4$, $F_{l}=3$; basis set Dyall-cvQZ+sp
Model $k$ [$10^{-10}$ Hz/((V/m)2)]
DCHF $-3.220$
SD9/10au $-3.375$
SD9/30au $-3.379$
SD9/1000au $-3.384$
SD8_SDT9/10au $-2.308$
SD8_SDT9/30au $-2.311$
SDT9/10au $-2.468$
SDT9/30au $-2.469$
SDT8_SDTQ9/10au $-2.112$
Exp.Simon_PRA1998 $-2.271(4)$
Exp.Godone_PRA2005 $-2.05(5)$
Safronova et al.beloy_derevianko_PRL2006 $-2.271(8)$
Angstmann et al.Angstmann_Dzuba_Flambaum_PRA2006 $-2.26(2)$
In view of the observed trends and their magnitude it is expected that the remaining small deviation from the most accurate experimental value for the Stark shift can further largely be removed by including full quadruple excitations in the wavefunction expansion. Such a calculation is not impossible but would be very time (and resource) consuming and beside the point. It has here been demonstrated that the presented method yields accurate hyperfine Stark shifts that systematically converge towards reliable reference results.
@@
With confidence in the method established in the aforegoing sections it is now applied to an atom where high-level reference results are not available and which has a more complex electronic structure. In this case the fractional occupation in the DCHF calculation is $f=15/16$ where $m=15$ represents the thirteen $4f$ electrons plus the two $6s$ electrons. This averaging was required to assure proper convergence of the DCHF wavefunction.
@@ III.4.1 ${}^{2}F_{7/2}(F=4)$${}^{2}F_{5/2}(F=3)$ clock transition
Denoting the static electric dipole polarizability as $\alpha_{J,M_{J}}$ the differential polarizability for the present levels is defined as $\Delta\alpha_{M_{J}}=\alpha_{\frac{5}{2},M_{J}}-\alpha_{\frac{7}{2},M_{J}}$. The relationship with the Stark coefficient can then be written as $-2k_{M_{J}}=\Delta\alpha_{M_{J}}$. The results in Table 3 are calculated using Eq. (6) and $\Delta\alpha_{M_{J}}$ is then derived from those values. In addition, individual static polarizabilities $\alpha_{J}$ are calculated by using the FF method. Generally, the single-field-point results match those from the more accurate FF method up to three significant digits when $E=0.0005$ a.u. is used. With $E=0.001$ a.u. correspondence is up to two digits.
Table 3: Stark frequency shift $k$ calculated according to Eq. (6), ground-state electric polarizability and differential polarizability for 169Tm ($4f_{7/2,5/2}$), $E_{z}=0.001$ a.u., $M_{J}=1/2$
Basis set/Model $k$ [$10^{-9}$ Hz/((V/m)2)] $\Delta\alpha$ [a.u.] $\alpha_{\frac{7}{2}}$ [a.u.]
QZ/DCHF $1.42$ $-0.115$ $186.6$
QZ+/DCHF $1.44$ $-0.115$
QZ/SD15_10au $0.52$ $-0.042$ $161.2$
QZ/SDT15_10au $1.77$ $-0.142$
QZ/SD23_10au $1.34$ $-0.108$ $151.2$
QZ/SD33_10au $1.57$ $-0.126$ $151.8$
QZ/SD41_10au $1.58$ $-0.127$
final $2.83$ $-0.23\pm 0.11$
Experiment $-0.063(30)$ NatComm_Tm_2019 $130\pm 16$ Ma_Tl_pol

### v2

@@ III.2.2 Hyperfine Stark Effect
For the electronic ground state corresponding to the valence configuration $4s^{1}$ the total electronic angular momentum quantum number is J=1/2{J=1/2}. The considered isotope has $I=3/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=2$ and $F_{l}=1$. The fractional occupation in the DCHF calculation is $f=1/2$.
@@
Results for the hyperfine Stark coefficient are shown in Table 2 for various electronic-structure models and are compared with experimental and theoretical literature results. As a general effect, the hyperfine constant of a given atom in an electronic $S$ state diminishes when an external $E$ field is included. This observation is explained by the fact that the $E$ field partially shifts spin density from $s$ wave to $p$ wave character in the atomic ground state (through $s$-$p$ mixing in the polarized atom), thus reducing the hyperfine interaction.
The hyperfine Stark coefficient in mean-field approximation (DCHF, both basis sets) differs from the experimental result by more than 2020% and lowest-order electron correlation effects from the $4s,4p,5s$ Rb shells even increase this discrepancy (model SD9, both basis sets). Upon including combined triple excitations from the $4s,4p$ shells and the $5s$ valence shell (model TZ/SD8_SDT9) a strong correction to $k$ is obtained. The inclusion of full triple excitations (model TZ/SDT9), i.e., including those from the $4s,4p$ shells, yields another large change in $k$. Close agreement with the result obtained by Safronova et al. Safronova_BBR_2009 is achieved at this level of calculation. This is not surprising since the wavefunction model used in Ref. Safronova_BBR_2009 for obtaining the cited result is a linearized coupled cluster expansion including up to (perturbative) valence triple excitations which is similar to the present SDT9 model.
Table 2: Stark hyperfine coefficient $k$ for 87Rb with nuclear magnetic moment $\mu=2.75131\mu_{N}$ ISOLDE_BWeffect_1993
Model $k$ $10^{-10}$ Hz/((V/m)2)]
TZ/DCHF 1.522-1.522
TZ/S8_SD9/10au 1.138-1.138
TZ/SD9/10au 1.618-1.618
TZ/SD8_SDT9/10au 1.220-1.220
TZ/SDT9/10au 1.273-1.273
TZ/SDT8_SDTQ9/10au 1.159-1.159
TZ/SDTQ9/10au 1.228-1.228
TZ/SDTQ8_SDTQQ9/10au 1.158-1.158
TZ/SDTQQ9/10au 1.161-1.161
TZ/SD18_SDT19/10au 1.253-1.253
TZ/SD26_SDT27/10au 1.295-1.295
TZ/SD34_SDT35/10au 1.299-1.299
TZ/SD36_SDT37/10au 1.301-1.301
QZ+/DCHF 1.546-1.546
QZ+/SD9/10au 1.662-1.662
QZ+/SD9/30au 1.669-1.669
QZ+/SD8_SDT9/30au 1.202-1.202
QZ+/SDT9/10au 1.265-1.265
QZ+/SDT9/30au 1.271-1.271
Final $\boldsymbol{-1.234\pm 0.0223}$}
Exp. Mowat_Li_1972 1.23(3)-1.23(3)
Safronova et al. Safronova_BBR_2009 1.272-1.272111preliminary value
Angstmann et al. Angstmann_Dzuba_Flambaum_PRA2006 1.24(1)-1.24(1)
It becomes clear that higher excitation ranks in the wavefunction expansion have a much greater impact on the hyperfine Stark coefficient than improvements in the atomic basis set. It is thus attempted to systematically converge $k$ with respect to the former effects. The evolution of the results for these systematically improved models is displayed graphically in Fig. 1 for ease of comparison.
Figure 1: Stark hyperfine coefficient $k$ for the electronic ground state of the rubidium atom using various electronic-structure models (black: present QZ+, blue: present TZ) and compared with other theoretical results and experiment, including the experimental uncertainty; for three of the models the evolution of the result with the number of CI iterations is also shown.
Along the improving model series SD9 – SD8_SDT9 – SDT9 – SDT8_SDTQ9 – SDTQ9 $\ldots$ results strongly oscillate, even after having included combined quadruple excitations. However, the partial series of corresponding models SD8_SDT9 – SDT8_SDTQ9 – SDTQ8_SDTQQ9 does lead to a sufficiently converged result. Likewise, the partial series SD9 – SDT9 – SDTQ9 – SDTQQ9 also reaches a sufficiently converged value that is not far from the result with the model SDTQ8_SDTQQ9. It can be inferred that the result from the model SDTQQ9 that includes full quintuple excitations for the $9$ outermost electrons is near the Full CI result in the TZ basis set.
The final value is obtained as follows. The result from the model SDTQQ9 serves as the base value and corrections from correlations of inner-shell electrons and from the atomic basis set are added. In detail,
$k({\text{final}})$ $=$ $k({\text{TZ/SDTQQ9/10au}})$ (21)
$+k({\text{TZ/SD36\_SDT37/10au}})-k({\text{TZ/SD8\_SDT9/10au}})$
$+k({\text{QZ+/SDT9/10au}})-k({\text{TZ/SDT9/10au}})$
This result differs from the experimental central value by only about 0.30.3% and is the theoretical result that comes closest to the experimental value. The uncertainty for the present final value is obtained by addition of individual uncertainties for interelectron correlation effects, the basis-set approximation and the approximation due to the use of the Dirac-Coulomb Hamiltonian operator (which neglects the Breit interaction and QED effects) for obtaining the atomic wavefunctions. The former two uncertainties are obtained from the difference between the most elaborate and the second most elaborate models, respectively. The latter uncertainty is estimated to be 1%.
@@
This important isotope of cesium has received considerable attention in the past. For the electronic ground state corresponding to the valence configuration $6s^{1}$ the total electronic angular momentum quantum number is J=1/2{J=1/2}. The considered isotope has $I=7/2$ and the resulting hyperfine quantum levels are denoted as $F_{u}=4$ and $F_{l}=3$. The fractional occupation in the DCHF calculation is $f=1/2$.
@@
Results for 133Cs are compiled in Table 3. As in the case of 87Rb the mean-field result for $k$ is too large on the absolute, compared with the cited reference values, by around 4040%. Again, the inclusion of lowest-order electron correlation effects from the outermost atomic shells even increases this deviation. It is here in addition and for the model QZ+/SD9 shown that the consideration of virtual spinors of very high energy (up to $1000$ a.u. ) in the wave-function expansion does not affect the results significantly. Also as for 87Rb combined triple excitations from the outermost shells yield a very important correction to $k$. Including full triple excitations (model QZ+/SDT9/30au) yields a result that deviates from the most accurate literature results by roughly +9+9%. Accounting for combined quadruple excitations quenches the deviation to 7-7% where again, as in the Rb atom, the correction slightly overshoots the exact value. This latter model comprises about $1.9$ billion ($10^{9}$) terms in the CI expansion when the QZ+ basis set is used.
@@
Table 3: Stark hyperfine frequency shift $k$ for 133Cs ($5s_{1/2}$), $I=3.5$, $E_{z}=0.001$ a.u., hyperfine level quantum numbers are $F_{u}=4$, $F_{l}=3$.
Model $k$ $10^{-10}$ Hz/((V/m)2)]
TZ/DCHF 3.164-3.164
TZ/SD9/10au 3.272-3.272
TZ/SD8_SDT9/10au 2.319-2.319
TZ/SDT9/10au 2.433-2.433
TZ/SDT8_SDTQ9/10au 2.153-2.153
TZ/SDTQ9/10au 2.321-2.321
TZ/SDTQ8_SDTQQ9/10au 2.156-2.156
TZ/SDTQQ9/10au 2.161-2.161
TZ/SD19/10au 3.335-3.335
TZ/SDT19/10au 2.516-2.516
TZ/SDT27/10au 2.595-2.595
TZ/SD27/10au 3.395-3.395
TZ/SD37/10au 3.396-3.396
TZ/SD45/10au 3.380-3.380
QZ+/DCHF 3.220-3.220
QZ+/SD9/10au 3.375-3.375
QZ+/SD9/30au 3.379-3.379
QZ+/SD9/1000au 3.384-3.384
QZ+/SD8_SDT9/10au 2.308-2.308
QZ+/SD8_SDT9/30au 2.311-2.311
QZ+/SDT9/10au 2.468-2.468
QZ+/SDT9/30au 2.469-2.469
QZ+/SDT8_SDTQ9/10au 2.112-2.112
Final $\boldsymbol{-2.347\pm 0.084}$}
Exp.Simon_PRA1998 2.271(4)-2.271(4)
Exp.Godone_PRA2005 2.05(5)-2.05(5)
Safronova et al.beloy_derevianko_PRL2006 2.271(8)-2.271(8)
Angstmann et al.Angstmann_Dzuba_Flambaum_PRA2006 2.26(2)-2.26(2)
Even higher excitation ranks can with the present method only be treated in the smaller TZ basis set. There, even a calculation with full quadruple excitations does not yet deliver a converged result (model TZ/SDTQ9). A converged result, however, is obtained when at least combined quintuple excitations are included in the wavefunction expansion.
Fig. 2 displays a selection of results supporting the present discussion. Correlations among and with the electrons from inner shells down to and including the $3s$ shell (model SD45) lead to a small increase of $k$, on the absolute. A very similar correction is obtained when full triple excitations are included in these expansions (models SDT9 and SDT27). This suggests that there is a negligible effect from including even higher excitation ranks in order to obtain corrections from inner-shell electron correlations.
Figure 2: Stark hyperfine coefficient $k$ for the electronic ground state of the cesium atom using various electronic-structure models (black: present QZ, blue: present TZ) and compared with other theoretical results and experiment, including the experimental uncertainty; for two of the models the evolution of the result with the number of CI iterations is also shown.
The final present value of the hyperfine Stark shift for Cs is, therefore, obtained as follows. The base value is taken from the model TZ/SDTQ8_SDTQQ9. This result is corrected by inner-shell correlations from the $4s,4p,4d$ shells at the SDT level and for inner-shell correlations from the $3s,3p,3d$ shells at the SD level. A small basis-set correction is added from the SDT9 model as well as a correction from including virtual spinors up to 1000 a.u. (model QZ+/SD9). Mathematically, the corresponding evaluations are
$k({\text{final}})$ $=$ $k({\text{TZ/SDTQ8\_SDTQQ9/10au}})$ (22)
$+k({\text{TZ/SDT27/10au}})-k({\text{TZ/SDT9/10au}})$
$+k({\text{TZ/SD45/10au}})-k({\text{TZ/SD27/10au}})$
$+k({\text{QZ+/SD9/1000au}})-k({\text{QZ+/SD9/10au}})$
$+k({\text{QZ+/SDT9/10au}})-k({\text{TZ/SDT9/10au}})$
The uncertainty on this final result is estimated in the same way as has been done for the above Rb atom.
The present final result is compatible with the measurement by Simon et al. from 1998 Simon_PRA1998 and incompatible with the measurement by Godone et al. from 2005 Godone_PRA2005 . The present result is also compatible with the theoretical results by Safronova et al. beloy_derevianko_PRL2006 and by Angstmann et al. Angstmann_Dzuba_Flambaum_PRA2006 . The result by Godone et al. is in conflict with all other experimental and theoretical results. The present result supports previous theoretical values as well as the measurement by Simon et al.
@@
With confidence in the method established in the aforegoing sections it is now applied to an atom where high-level theoretical reference results are not available and which has a more complex electronic structure. In this case the fractional occupation in the DCHF calculation is $f=15/16$ where $m=15$ represents the thirteen $4f$ electrons plus the two $6s$ electrons. This averaging was required to assure proper convergence of the DCHF wavefunction.
@@ III.4.1 $^{2}F_{7/2}(F=4)$$^{2}F_{5/2}(F=3)$ clock transition
Individual static polarizabilities $\alpha_{J}$ are calculated by using the FF method and are given in Table 4. Using Eq. (16) the $M_{J}$-dependent values of $\alpha_{D}$ are in the case of thulium states with $J=7/2$ obtained from the inverted equations
$\alpha_{0}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{6}\alpha_{D}\left({}^{2}F_{7/2,5/2}\right)-\frac{1}{6}\alpha_{D}\left({}^{2}F_{7/2,7/2}\right)$ (23)
$\alpha_{2}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{6}\left[\alpha_{D}\left({}^{2}F_{7/2,7/2}\right)-\alpha_{D}\left({}^{2}F_{7/2,5/2}\right)\right]$ (24)
$\alpha_{0}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{5}{2}\alpha_{D}\left({}^{2}F_{7/2,3/2}\right)-\frac{3}{2}\alpha_{D}\left({}^{2}F_{7/2,1/2}\right)$ (25)
$\alpha_{2}\left({}^{2}F_{7/2}\right)$ $=$ $\frac{7}{2}\left[\alpha_{D}\left({}^{2}F_{7/2,3/2}\right)-\alpha_{D}\left({}^{2}F_{7/2,1/2}\right)\right]$ (26)
results for which are also given in Table 4.
Table 4: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}F_{7/2}$ multiplet
$\alpha_{D}$
Model $^{2}F_{7/2,1/2}$ $^{2}F_{7/2,3/2}$ $^{2}F_{7/2,5/2}$ $^{2}F_{7/2,7/2}$ $\alpha_{0}(^{2}F_{7/2})$ $\alpha_{2}(^{2}F_{7/2})$ calc. from $\{M_{J}\}$
TZ/DCHF $186.4554$ $185.7002$ $184.5674$ 2.643-2.643 $1/2,3/2$
TZ/SD15/6au $163.9967$ $163.2578$ $162.1495$ 2.586-2.586 $1/2,3/2$
TZ/SD13_SD15/6au $163.9897$ $163.2515$ $162.1442$ 2.584-2.584 $1/2,3/2$
TZ/SD13_SDT15/6au $161.9754$ $161.3215$ $160.3407$ 2.289-2.289 $1/2,3/2$
TZ/SD13_SDTsppdQ15/6au $167.2591$ $166.5212$ $165.4144$ 2.583-2.583 $1/2,3/2$
QZ/DCHF $186.6210$ $185.8505$ $184.6948$ 2.697-2.697 $1/2,3/2$
QZ/SD15/10au $161.2430$ $160.4891$ $158.9842$ $156.7269$ $159.3583$ 2.639-2.639 $1/2,3/2$
$159.3604$ 2.634-2.634 $5/2,7/2$
QZ/SDT15/10au $160.8884$ $160.2111$ $159.1952$ 2.371-2.371 $1/2,3/2$
QZ/SD23/10au $151.2078$ $150.4219$ $149.2431$ 2.751-2.751 $1/2,3/2$
QZ/SD33/10au $151.8781$ $151.0859$ $149.5087$ $147.1326$ $149.8976$ 2.773-2.773 $1/2,3/2$
$149.9047$ 2.772-2.772 $5/2,7/2$
Experiment $130\pm 16$ Ma_Tl_pol